OpenStudy (rissa):

Solve the inequality. -5.3 >= 6.7 + 4.3 + q

OpenStudy (anonymous):

-16.3>=q

OpenStudy (anonymous):

-16.3>=q

OpenStudy (anonymous):

\[-16.3\ge q\]

OpenStudy (anonymous):

-5.3 >= 6.7 + 4.3 + q steps: -5.3-6.7-4.3 >= q -16.3 >= q